OpenStudy (anonymous):

(x^2-k)(x+k)=x^3+k(x^2-x-5) What is the value of k?(K isnt 0)

OpenStudy (anonymous):

x^3 - kx + kx^2 - k^2 = x^3 + kx^2 - kx - 5k -k^2 = - 5k 5k = k^2 k = 5

OpenStudy (anonymous):

Thank you jimmyrep

OpenStudy (anonymous):

you r welcome